OpenStudy (anonymous):

solve each equation and figure out what x equals to : x/4-3=5 x=? (6x-2)/8=2 x=? (5x+3)/4+5=17 x=? (7x+1/9)+11=15 x=?

OpenStudy (anonymous):

x/4-3 = 5 move -3 to the right hand side and change its sign x/4 = 8 times both sides by 4 x = 32 do the rest similiarly o-o.., keep making changes to both sides to simplify the equation

OpenStudy (anonymous):

move the numbers to the other side of x to solve it (x/4) - 3 = 5 x/4 = 8 x =32 (6x - 2)/8 = 2 6x - 2 = 16 x = 3 (5x + 3)/4 + 5 = 17 5x + 3 = 48 x = 9 (7x + 1)/9 + 11 = 15 (7x + 1)/9 = 4 (7x +1) = 36 x = 5

OpenStudy (anonymous):

Thanks!!!!!!! :)

OpenStudy (anonymous):

sure ^^

OpenStudy (anonymous):

That was really helpful :)

OpenStudy (anonymous):

Your good at math!

OpenStudy (anonymous):

glad you feel that way hahaha actually not really ^^; thx for the medal (Y)

OpenStudy (anonymous):

Your welcome:)